OpenStudy (anonymous):

Simplify by removing factors of 1. 9p^2+9p/45p^2+81p

OpenStudy (anonymous):

\[\frac{9p^2+9p}{45p^2+81p}\] \[\frac{9p(p+1)}{9p(5p+9)}\]

OpenStudy (anonymous):

cancel, get \[\frac{p+1}{5p+9}\]

OpenStudy (anonymous):

Thanks!!! :-)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Can you please do one more for me?

OpenStudy (anonymous):

Whacha got?

OpenStudy (anonymous):

post you will get lots of answers

OpenStudy (anonymous):

Factor completely. 6x^5-48x^4+12^3

OpenStudy (anonymous):

alisa, I think you probably made a typographical error on the question. I think the expression is probably\[6x^5-48x^4+12x^3=6x^3(x^2-8x+2)\]The trinomial factor has irrational roots that don't seem to factor with integer coefficients.